OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -3x-10y=7

OpenStudy (anonymous):

y=-3/10x-7/10 slope -3/10 y intercept -7/10

OpenStudy (anonymous):

\[-3x - 10y = 7\] To get the equation in slope intercept form, you have to solve for y \[-10y = 7+3x\] \[y = \frac{ 7 }{ -10 } + \frac{ 3x }{ -10 }\] The number next to the x is the slope and the number without the x is your y-intercept. Therefore, -3/10 is your slope and your y-intercept is -7/10

OpenStudy (anonymous):

Thank you both so much!